OpenStudy (anonymous):

what does the graph of r = theta, theta >= 0 looks like?

OpenStudy (anonymous):

you cant graph on this website

OpenStudy (anonymous):

\[\theta=n \pi, n \in Z:intersection.with.polar.axis\] \[(1/2)n \pi, n \in odd Z: intersection.with.(1/2) \pi.axis.\] \[when.r=0, \theta=0\] counterclockwise spiral.

OpenStudy (anonymous):

thanks :)) nice :))

myininaya (myininaya):

\[r^2=x^2+y^2, x=r \cos(\theta), y=r \sin(\theta)\] \[x=r \cos(\theta)=> \frac{x}{r}=\cos(\theta)=> \cos^{-1}(\frac{x}{r})=\theta\] \[r=\theta=>\sqrt{x^2+y^2}=\cos^{-1}(\frac{x}{r})=>\sqrt{x^2+y^2}=\cos^{-1}(\frac{x}{\sqrt{x^2+y^2}})\] \[x^2+y^2=[\cos^{-1}(\frac{x}{\sqrt{x^2+y^2}})]^2\] and if for some reason you also want to write as a Cartesian equation here you go

OpenStudy (anonymous):

thanks guys :))